Find the maximum net charge that can be placed on a spherical conductor of radius 64 cm before dielectric breakdown of the air occurs. The dielectric strength of the air is 3
Answer #1

The electric field strength on the surface of the sphere E = k * Q/ R^2

Q = E * R^2 / k


= 3 * 10^6 * 0.4096 / (8.99 * 10^9)


= 0.1367 * 10-3 C


= 136.7 muC


----------------------------------------------------------

Potential f = E * R


= 3 * 10^6 * 64 * 10-2


= 1920000 V


= 1920 kV